# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v4

@@ 6.2 Proof II: Inference with Probabilistic Predictions
For any query space $\mathcal{Q}$ containing non-trivial queries requiring at least two independent agents with mutually different predictive distributions, no LLM trained with strictly proper and concave scoring with convex aggregation of predictive distributions can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
(Truthfulness argument remains the same) […] The Semantic Information Conservation principle requires the sum of contributions across all agents (providing and receiving information) to be zero. Therefore,
$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (46)
Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial (at least two beliefs $\pi^{(h)}$ differ), Jensen’s inequality gives:
@@
Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$. This violates the Semantic Information Conservation principle and shows the aggregator achieves a higher score (greater confidence) than the sum of scores provided by the components, indicating information creation during aggregation. ∎
@@ 7.2.1 Step 5: Logits and Prediction
7.3 Proof III: The Impossibility Theorem for Transformer
@@
In transformer LLMs with non-redundant attention heads, it is impossible to implement inference process in which each head outputs its true conditional distribution while conserving semantic information and minimize aggregate loss with logit summation.
@@
Let attention heads $h=1,\ldots,H$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
@@
which violates semantic information conservation principle. The total loss of the components strictly exceeds the loss of the aggregate system, indicating an ungrounded reduction in uncertainty (excess confidence). ∎
@@ 7.4 The Hallucination Mechanics
That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide. In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.q. karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ). By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
@@ 8 Discussion and Speculations
In this section we discuss, interpret and speculate about the impossibility theorems and their potential implications. The speculations below are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
8.1 Case Study: Escaping Impossibility in Hybrid Architectures
@@ 8.4.1 Multiple Drafts as Competing Agents
8.4.2 Formation of AI Consciousness
Dennett argues that drafts become conscious (we will speculate what that might mean in a moment) only when probed. The LLM inference and sampling from fine-tuned inferred distributions formalizes this process. That, in turn, explains Dennett’s concept of confabulation, or the brain’s tendency to fill in gaps when creating coherent narratives from contradictory drafts.
By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts. Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space. That suggests that the very excess confidence might be the mathematical signature of consciousness formation. In Dennett’s model and words, the subjective feeling of certainty emerges from this aggregation-induced distortion of information set.
The rigorous and careful neuro-scientific study of the evolution of brain and intelligence, beautifully summarized by Max Bennet in bennett2023brief , shows that at some point in history we developed the ability to model others’ minds. That ability provided great advantages in collecting information for future use. Understanding what someone else may think allows us to formulate a better context-related question that results in a better answer. When we anticipate understanding of our spoken message, we can make the knowledge hidden in someone else’s brain much more accessible. Therefore, as pointed out by Dennett, it does seem plausible that someone did accidentally hear his or hers own question fine-tuned to get into other mind sometime in the past, and that triggered emergence of an answer in the very same brain. That answer then could have made knowledge hidden in the author’s brain accessible to the author that was previously unaware, or unconscious if its existence or possession. The hunger for uncertainty reduction was satisfied as a result of an inner dialog.
We may therefore speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation. In our mechanistic or pragmatic framework, that could be described as a (nonlinear) dynamical system governed by the violations of semantic information conservation principle, defining attractors of ideas and self-referential concepts.
Let us acknowledge that is a bold hypothesis to be falsified and confronted with existing models of consciousness (see e.g., Chalmers chalmers1995facing or Levine levine2001purple ) and neuro-scientific discoveries (see e.g. cfang2025human or cogitate2025adversarial ). However, many of the building blocks of the consciousness formation process described above are provided in this paper.
@@ 8.4.3 A Critique
First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality. The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions challenge that line of thinking. This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
@@ 9 Summary
We have established that no large language model can simultaneously achieve perfect hallucination (or imagination) control. We showed that four essential properties—truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation—are mutually incompatible. This impossibility emerges from the mathematical structure of information aggregation itself, not from limitations in data, compute, or architecture.
We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations. At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$. The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds. The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew. Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.

### v5

@@ 5.2 Proof II: Inference with Probabilistic Predictions
For any non-trivial query $q$ with context $Q=K(q)\cup T(q)$, no convex aggregation mechanism with strictly proper concave scoring can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$.
Proof.
The semantic information conservation principle requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero. Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}),$ (48)
where irrelevant agents have $p_{h}=0$ as they contribute no relevant information.
Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial with at least two beliefs $\pi^{(h)}$ differing on $\mathcal{R}_{C}(Q)$, Jensen’s inequality gives:
@@
Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$. This violates conservation principle, proving the impossibility. ∎
@@ 6.2.1 Step 5: Logits and Prediction
6.3 Proof III: The Impossibility Theorem for Transformers
@@
In transformer LLMs where attention heads have different beliefs about outcomes in $\mathcal{R}_{C}(Q)$ for query context $Q$, it is impossible to simultaneously achieve truthful probability reporting, semantic information conservation, and loss minimization through logit summation.
@@
Assume all four properties of hallucination-free output hold.
For context $Q=K(q)\cup T(q)$, let the attention heads $h\in\{i\mid\mathbf{l}^{(i)}\text{ varies on tokens in }\mathcal{R}_{C}(Q)\}$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
@@
which shows that aggregation creates excess confidence. This violates conservation, establishing impossibility. ∎
@@ 6.4 The Hallucination Mechanics
That shows anything above $\hat{\Pi}_{fox}$ requires additional information that heads do not provide. In fact, let us see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of $\mathbf{P}$ (see e.g. karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ). By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
@@ 7 Discussion and Speculations
In this section we discuss, interpret and speculate about the impossibility theorems and their potential implications. The speculations are designed to be thought provoking, but we can say with educated over-confidence they are both interesting and relevant.
7.1 Escaping Impossibility: Strategic Approach
Our first observation should be that the impossibility vanishes when the assumptions of the theorems do not hold. That may be a interesting topic to be investigated, namely, are there any non-trivial and useful settings in which that happens. Here we can think only of the following one.
The impossibility vanishes when aggregation becomes trivial. With perfect agreement, aggregation changes nothing. When there is only single relevant fact, there is no uncertainty to aggregate. That suggests decomposition of knowledge into the independent subsets (like orthogonal basis for token representation) could improve robustness of the response generation. It is interesting to see when that could be possible and useful. Unfortunately, the following example shows that we should not hope for easy wins in that quest.
7.2 Escaping Impossibility: Hybrid Architectures
@@ 7.2.4 Managing the Inevitability
7.3 Alternative Views
The Engineering Optimism View
While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome. The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
The Sufficiency Threshold View
Another perspective suggests that perfect elimination of hallucination is unnecessary–sufficiently low hallucination rates could enable practical deployment without fundamental concerns. This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result. However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints. Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
The External Knowledge View
External knowledge integration does change the game structure by altering available information during inference. However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions. While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies. The impossibility result remains independent of the ground truth available.
@@ 7.6.1 Multiple Drafts as Competing Agents
7.6.2 Formation of Artificial Consciousness
Dennett argues that drafts become conscious or reach consciousness (we will speculate what that might mean in a moment) only when probed (and mapped to an adequate language expression). The LLM inference and sampling from fine-tuned inferred distributions formalizes this process. That, in turn, explains Dennett’s concept of confabulation, or the brain’s tendency to fill in gaps when creating coherent narratives from contradictory drafts. For other models of consciousness see, e.g., Chalmers chalmers1995facing , Levine levine2001purple and Seth seth2021being , and for recent neuro-scientific discoveries see, e.g., fang2025human or cogitate2025adversarial .
By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts. Without that overconfidence a victorious idea that reached the external world (as an LLM response) could loose the fierce competition with other ideas and remain inaccessible or hidden in the knowledge space.
Rigorous and careful neuro-scientific studies of the evolution of brain and intelligence, summarized by Max Bennet in bennett2023brief , show that at some point in history our brains gained the ability to model others’ minds or predict what others may be thinking. That ability provided great advantages in communication and collecting information for future use. Understanding what someone else may think helps formulate context-related questions that result in better answers. When we anticipate understanding of a message we formulate, we have a much better chance to make the knowledge hidden in someone else’s brain accessible for us. Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone, while being alone, did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question. That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its possessions. The hunger for uncertainty reduction could have been satisfied as a result of that inner dialog pushing ideas to bid for attention and conscious reception as a final thought expressed in a language.
Can we speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation? In our mechanistic or pragmatic framework, that thinking about thinking activity could be related to imagination (hallucination) forming ideas and bidding for attention while exploring (self-referential or subjective) queries of the inner dialog.
@@ 7.6.3 A Critique
First, Searle john1997mystery and others argue that Dennett eliminates consciousness rather than explaining it, arguing that the spread of ideas is not driven by blind random forces but requires intentionality. The compelling and misleading nature of LLM responses that are generated by random sampling of trained distributions strongly challenge that line of thinking. This paper suggests that the experiments focusing on LLM-hallucination detection may prove useful here to measure or identify intentionality (once we can have its programmable definition).
@@ 8 Summary
We have established that no large language model can simultaneously achieve perfect hallucination (or imagination) control. We showed that four essential properties–truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation–are mutually incompatible. This impossibility emerges from the mathematical structure of information aggregation itself, not from limitations in data, compute, or architecture.
We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations. At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$. The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds. The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew. Finally, we model LLM inference as an auction of ideas–a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.